OpenStudy (anonymous):

A certain tribe is distinguished by the fact that 45% of the males have 6 toes on their right foot.Determine using suitable approximation that the probability that in a group of 200 males from the tribe, more than 97 have six toes on their right foot.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

there is no standard deviation given

OpenStudy (anonymous):

that is the question...:S

OpenStudy (anonymous):

this is a binomial problem , either you have 6 toes (success) or you do not

OpenStudy (anonymous):

i think you use binomial but im not sure how to do it

OpenStudy (anonymous):

you can approximate the binomial using a normal distribution

OpenStudy (anonymous):

like?

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

you can also do this with calculators

OpenStudy (anonymous):

makes sense?

OpenStudy (anonymous):

so i still use the x- mean / standard deviation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get the first part and i got the n and the probability but im not sure how to continue. could you please show me the working if you get it?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

which part did you get

OpenStudy (anonymous):

actually my answer is very close but not quite. i did 97.7-90/ 7.03

OpenStudy (anonymous):

sorry 97.5-90/7.03

OpenStudy (anonymous):

mean = n*p std = sqrt( n*p*(1-p))

OpenStudy (anonymous):

hm i think i got the answer actually. i didnt do it by calculator so it won't be accurate.

OpenStudy (anonymous):

so if its more than 97 it should be 97.5?

OpenStudy (anonymous):

we want greater than 97, so we use P( x>= 97.5)

OpenStudy (anonymous):

so 97 to 97.5?

OpenStudy (anonymous):

yes, because the binomial is a rectangle , but we want smooth

OpenStudy (anonymous):

okay thank you so much for your help..:)

OpenStudy (anonymous):